Robinson's preferences between apples (a) and bananas (b) are expressed by the following:

 

U = (a+2)0.5(b+1)0.5

 

(a) Show that Robinson's indifference curves are negatively sloped.

(b) Are they convex to the origin? Explain.


1
Expert's answer
2021-12-06T16:37:36-0500

Robinson's interference curves


a.) From U = U = (a + 2)0.5 ∗(b+ 1)0.5

Totally differentiating the utility function

"du=\frac{\delta}{\delta a}[(a+2)^{0.5}(b+1)^{0.5}da]+\frac{\delta}{\delta b}[(a+2)^{0.5}(b+1)^{0.5}db]"


"du=0.5(a+2)^{-0.5}(b+1)^{0.5}da+0.5(a+2)^{0.5}(b+1)^{-0.5}db"


along an indifference curve change in utility du=0

Also we may apply (a) along horizontal axis hence,

"0.5(a+2)^{-0.5}(b+1)^{0.5}da+0.5(a+2)^{0.5}(b+1)^{-0.5}db=0"


"\frac{da}{db}=\frac{0.5(a+2)^{0.5}(b+1)^{-0.5}}{0.5(a+2)^{-0.5}(b+1)^{0.5}}=\frac{a+2}{b+1}"


"\frac{da}{db}= slope \ of\ indifference"


"=MRS_{a,b }" [Nominal rate of substitution between a and b]

As, "MRS_{a,b } \lt0" implies robinsons indifference curve is negatively sloped







b.) Now "|MRS_{a,b }|=\frac{a+2}{b+1}=\frac{a}{b+1}+\frac{2}{b+1}"


Now "\frac{\delta |MRS_{a,b }|}{\delta a}=\frac{1}{b+1}\gt0 \ as\ b\gt0"

hence they are convex to the origin